1. Find the alpha with LI
2. confirm if the vector is true
Encuentre el valor de a para que los siguientes vectores se sean LI. (2, -3,1),(-4,-6,1),(a, 1, 2) Cargar archivo
Compruebe si la igualdad es verdadera. [H] = R2; H = {(1,2),(-1,0)} Ť Cargar archivo
Answer #1

linearly independent oenot 2) Since H= To check His lil {(1,2), (-3,0) d (112) + B (-1,0)= (0,0) .:(2-13, 20) = (0:0) 2-3 = =Given vectors are, (21-3, d),(-416,2), (2,1,2) - He know that, the vectors are linearly independent if theit determinant is L